# The Unbounded Denominators Conjecture

### v1

@@ 5.1. Normalizations of $\widetilde{F}_{N}$
From the formal identity $\varphi_{N}(x)=\psi_{N}(x^{1/N})^{N}$, we deduce
$\varphi_{N}(z)=\gamma^{-N}_{N}\eta_{2}(z^{1/N})^{N}/\eta_{1}(z^{1/N})^{N}.$
Letting $\phi_{i}=\eta_{i}(z^{1/N})$, we deduce the lemma from an elementary manipulation directly from equation (5.0.3). ∎
5.2. The geometry of $\widetilde{\Gamma}_{N}$ and $\widetilde{\Phi}_{N}$
Recall that $\widetilde{F}_{N}(\tau|r_{N})=\zeta\widetilde{F}_{N}(\tau)$, where $r_{N}\in\mathrm{SO}_{2}(\mathbf{R})$ corresponds in $D(0,1)$ to multiplication by $\zeta$. Explicitly, we have
(5.2.1) $r_{N}=\left(\begin{matrix}\cos(\pi/N)&-\sin(\pi/N)\\ \sin(\pi/N)&\cos(\pi/N)\end{matrix}\right).$
The element $r_{N}$ acts transitively on the cusps which include $i\infty$, and hence $N$ of the cusps are given by $\cot(\pi k/N)$ for any integer $k$. For example, if $N$ is even, then we can take $k=(N/2)$ and see that 0 is a cusp.
Definition 5.2.2.
let $\widetilde{\Phi}_{N}$ denote the group $\langle\widetilde{\Gamma}_{N},r_{N}\rangle$.
Since $r_{N}$ normalizes $\widetilde{\Gamma}_{N}$, this contains $\widetilde{\Gamma}_{N}$ as a normal subgroup with $\Phi_{N}/\widetilde{\Gamma}_{N}\simeq\mathbf{Z}/N\mathbf{Z}$.
Theorem 5.2.3.
We have $\widetilde{\Phi}_{N}\simeq\mathbf{Z}/N\mathbf{Z}\ast\mathbf{Z}$ where $r_{N}$ is as in equation (5.2.1), and the $\mathbf{Z}$ is the stabilizer of $i\infty$ in $\widetilde{\Gamma}_{N}$ which is generated by
$t_{N}:=\left(\begin{matrix}1&2\cot(\pi/2N)\\ 0&1\end{matrix}\right).$
The group $\widetilde{\Gamma}_{N}$ is the free group on $N$ generators generated by $t=t_{N}$ and its conjugates by powers of $r=r_{N}$.
Proof.
The group $\widetilde{\Gamma}_{N}$ is certainly generated by the stabilizers of the cusps $c$ with $\widetilde{F}_{N}(c)\in\mu_{N}$. If we denote the generator of the stabilizer of $i\infty$ by
(5.2.4) $t:=\left(\begin{matrix}1&c_{N}\\ 0&1\end{matrix}\right),$
then the stabilizers of the other cusps associated to $\mu_{N}$ are generated by the conjugates of $t$ by $r$. But we know that $\widetilde{\Gamma}_{N}$ has finite covolume and is not compact. Consider the Dirichlet domain $\Omega_{N}$ associated to $\widetilde{\Gamma}_{N}$ around $z=0$ in the Poincaré model. We can describe $\Omega_{N}$ as the region:
$\Omega:=z\in\mathbf{C}\ \text{such that}\ d(gz,i)\geq d(z,i),\quad g,g^{-1}\in\{r^{k}tr^{-k}\},k=0,1,\ldots,N-1.$
The region such that $d(r^{k}tr^{-k}z,i)\geq d(z,i)$ and $d(r^{k}t^{-1}r^{-k}z,i)\geq d(z,i)$ in the Poincaré disc model is the region bounded by two geodesics starting at $\zeta^{k}$ going in opposite directions and intersecting the boundary at $\zeta^{k}e^{\pm i\theta}$ where $c_{N}=2\cot(\theta/2)$. There are exactly $2N$ such arcs corresponding to the $N$ generators and their inverses. In particular, if $\theta<\pi/N$ is too small, the fundamental region will have infinite volume, whereas if $\theta>\pi/N$ is too big, then the Dirichlet domain will only contain at most $N$ cusps, and yet $\mathbf{H}/\widetilde{\Gamma}_{N}$ has $N+1$ cusps. Thus we must have $c_{N}=2\cot(\pi/2N)$. ∎
The region $\Omega_{N}$ has $2N$ cusps given by half-integer powers of $\zeta=\exp(2\pi i/N)$. It follows from this that the cusp width at this remaining cusp will be $Nc_{N}=2N\cot(\pi/2N)$, and that all the remaining $N$ cusps are in the same orbit of $\widetilde{\Gamma}_{N}$. As an example, note that $e^{k\pi i/N}$ in the Poincaré disc model corresponds to
(5.2.5) $i\cdot\frac{1+e^{-k\pi i/N}}{1-e^{-k\pi i/N}}=\cot(\pi k/2N)$
in $\partial\mathbf{H}$. But now we have (for example)
(5.2.6) $r^{m}_{N}t_{N}r^{-m}_{N}\cdot\cot\left(\frac{\pi(2m-1)}{2N}\right)=\cot\left(\frac{\pi(2m+1)}{2N}\right).$
A fundamental domain for $\widetilde{\Phi}_{N}$ is given by the hyperbolic quadrilateral with vertices $0,\zeta^{-1/2},1,\zeta^{1/2}$. Translated to $\mathbf{H}$ this is bounded by geodesics from $i$ to $\cot(\pi/2N)$ to $i\infty$ to $\cot(\pi/2N)$ and back to $i$.
5.3. The group $\widetilde{\Psi}_{N}$
The group $\widetilde{\Phi}_{N}=\langle r_{N},t_{N}\rangle$ is contained with index two in the larger group $\widetilde{\Psi}_{N}=\langle s,t_{N}\rangle$ where $s^{2}=r_{N}$ is a rotation of order $2N$. The function $\widetilde{F}^{N}_{N}$ is invariant under $\widetilde{\Phi}_{N}$ taking the value $1$ at one cusp and $\infty$ at the other. From this we deduce:
Lemma 5.3.1.
There is an equality
(5.3.2) $1-\widetilde{F}^{N}_{N}(s\cdot\tau)=\frac{1}{1-\widetilde{F}^{N}_{N}(\tau)}.$
Proof.
Both sides of the equation are uniformizers of $\mathbf{H}/\widetilde{\Phi}_{N}$ which take the value $1$ at one cusp and $\infty$ at the other. This specifies them uniquely up to $z\mapsto\lambda z$ scalings. However, this last ambiguity is removed by noting that both sides have their (unique) zero at $\tau=i$. ∎
We also deduce from this the equation
$1-G_{N}(-q)=\frac{1}{1-G_{N}(q)}.$
The group $\widetilde{\Psi}_{N}$ has a fundamental domain consisting of the points $0,1,\zeta^{1/2}$. But this is none other than a hyperbolic triangle with angles $\{\alpha,\beta,\gamma\}=\{\pi/N,0,0\}$. But this suggests that $\widetilde{F}_{N}$ should directly be related to Schwarz triangle functions, which leads to a direct description of the inverse functions $\psi_{N}$ and $\varphi_{N}$ in terms of hypergeometric functions.
Definition 5.3.3.
Let $s_{N}(z)$ denote the function
@@ 5.3. The group $\widetilde{\Psi}_{N}$
Lemma 5.3.5.
The solutions $\phi_{1}$ and $\phi_{2}$ of the ODE of equation (5.1.6):
are given explicitly by
@@
In particular, we have $\varphi_{N}(z)=(\gamma^{-1}_{N}\phi_{2}(z)/\phi_{1}(z))^{N}=(\gamma^{-1}_{N}s_{N}(z))^{N}$ and $\psi_{N}=\gamma^{-1}_{N}s_{N}(z^{N}).$
Proof.
This is elementary, since one can directly check that both sides satisfy the same differential equation then check the leading terms. ∎
Theorem 5.3.8.
The conformal radii of $F_{N}:D(0,1)\rightarrow\mathbf{C}\smallsetminus\mu_{N}$ and $G_{N}:D(0,1)\rightarrow\mathbf{C}\smallsetminus\{1\}$ are equal to $\gamma_{N}$ and $\gamma^{N}_{N}$ respectively, where
(5.3.9) $\gamma_{N}=\frac{\displaystyle{\Gamma\left(\frac{N-1}{2N}\right)^{2}\Gamma\left(1+\frac{1}{N}\right)}}{\displaystyle{\Gamma\left(\frac{N+1}{2N}\right)^{2}\Gamma\left(1-\frac{1}{N}\right)}}.$
We have a (convergent) expansion for $\gamma_{N}$ as follows:
(5.3.10) $\gamma_{N}=16^{1/N}\left(1+\frac{\zeta(3)}{2N^{3}}+\frac{3\zeta(5)}{8N^{5}}+\cdots\right).$
Proof.
From equation (5.1.3), it suffices to compute the limit of $\phi_{2}/\phi_{1}$ as $z\rightarrow 1$, which we can do directly given the explicit form in terms of hypergeometric functions (see also (5.4.3)). Similarly, the expansion can be derived directly from the expression in terms of the Gamma function. More precisely, we can write
(5.3.11) $\log\gamma_{N}=\frac{\log 16}{N}+\sum_{k=1}^{\infty}\frac{(2^{2k}-1)}{2^{2k-1}(2k+1)}\cdot\frac{\zeta(2k+1)}{N^{2k+1}}.$
@@
5.4. Uniform asymptotics for hypergeometric functions
@@
Fix a real constant $M_{0}>0$. Let $\psi$ be the digamma function. For $M\geq M_{0}$ and $|z|<e^{-MN}$, we have the uniform estimate:
(5.4.2) $\left|\frac{s_{N}(1-z)}{\gamma_{N}}-(1-z)^{1/N}\frac{\log z+2\gamma-2\psi(1/2+1/2N)}{\log z+2\gamma-2\psi(1/2-1/2N)}\right|\ll\frac{|z|}{N}$
where the implied estimates depend on $M_{0}$ but not on $N,M$.
@@
Let $a\notin\mathbf{Z}$. For $|z|<1$, we have the following equality:
@@
Note that the hypergeometric function is multivalued around $z=1$, but all different branches are accounted for by the branches of the logarithm. The function $s_{N}(1-z)$ is given by the ratio (up to a factor of $(1-z)^{1/N}$) by these functions for $a=1/2\pm 1/2N$. As $a\rightarrow 1/2$, the coefficients in this power series are uniformly bounded. For example, since $|a|<1$ we have $|(a)_{k}|/k!<1$, whereas $|\psi(k)-\psi(k+a)|$ is maximized when $k=1$ and a=1/21/4a=1/2-1/4. This immediately leads to the uniform estimates:
$\left|\sum_{k=1}^{\infty}\frac{(a)_{k}(a)_{k}}{k!^{2}}z^{k}\right|,\quad\left|\sum_{k=1}^{\infty}\frac{(a)_{k}(a)_{k}}{k!^{2}}z^{k}\left(2(\psi(k+1)-\psi(k+a))\right)\right|\ll|z|\leq e^{-MN}$
For $|z|<e^{-MN}$, we also have $\mathrm{Re}(\log z)\leq-MN$, and so in particular $|\log z|\geq MN$ regardless of the branch of logarithm. This leads to the estimate
$\frac{s_{N}(1-z)}{\gamma_{N}(1-z)^{1/N}}=\frac{\log(z)+2\psi(1)-2\psi(1/2+1/2N)+O(z)}{\log(z)+2\psi(1)-2\psi(1/2-1/2N)+O(z)}$
@@ 5.5. The region $F^{N}_{N}\sim 1$ and $F_{N}\sim\infty$
Fix a real number $M_{0}>0$. Consider $M\geq M_{0}$ and let $\Omega_{N}$ be a fundamental domain for $\Phi_{N}$. If
$\|F_{N}(\tau)^{N}-1\|<e^{-MN}$
for $\tau\in\Omega_{N}$, then for any $\epsilon>0$
(5.5.2) $\mathrm{im}(\tau)>\frac{2N^{2}M}{\pi^{2}}(1-\varepsilon)$
for $N\gg 1$, where the implied constant depends only on $M_{0}$ and $\epsilon$.
@@
Note that $F^{N}_{N}$ is $1:1$ on $\Omega_{N}$ and the only cusp where $F_{N}=1$ is at $\infty$, so it suffices to consider $F^{N}_{N}$ in a neighbourhood of the cusp $i\infty$. We may also take the branch of $F_{N}$ on this domain so that $F_{N}(i\infty)=1$. For $N$ sufficiently large, the inequality $\|F_{N}(\tau)^{N}-1\|<e^{-MN}$ implies that $|F_{N}(\tau)-1|<e^{-M(1-\varepsilon)N}$ for some $\varepsilon$ that tends to zero as $N$ increases. Hence, replacing $M$ by $M(1-\varepsilon)$, we may equivalently assume that, for the branch of $\psi_{N}$ such that $\psi_{N}(1)=1$ and $\widetilde{\psi}_{N}(1)=i\infty$, that
$\tau=\widetilde{\psi}_{N}(1-z),\qquad|z|<e^{-MN}.$
@@
Using once more the estimate $(1-z)^{N}\sim 1-Z$ where $Z<e^{M(1+\varepsilon)N}$ for sufficiently large $N$, we may consider
(5.5.4) $\tau=i\cdot\frac{\gamma_{N}+s_{N}(1-Z)}{\gamma_{N}-s_{N}(1-Z)},\quad|Z|<e^{-MN}.$
Now by Lemma 5.4.1 the leading term of this is
$\tau\sim-\frac{i\cot(\pi/2N)}{\pi}\left(2\gamma+\log x-\psi(1/2-1/2N)-\psi(1/2+1/2N)\right)$
The imaginary part of this does not depend on the choice of branch of $\log x$ and indeed only depends on $|x|$, and we deduce with this approximation that
$\mathrm{im}(\tau)\geq-\frac{\cot(\pi/2N)}{\pi}\left(2\gamma-NM+\psi(1/2-1/2N)+\psi(1/2+1/2N)\right)\sim\frac{2N^{2}M}{\pi^{2}}$
as required. ∎
We use this to derive a coarse upper bound on $\sup_{|z|=r}\log|F_{N}|$, which is not optimal but is enough as an input for the logarithmic error term in the Nevanlinna theory estimate in §6.
A horoball in the Poincaré disc model is a ball bounded by a circle inside the disc $D(0,1)$ which is tangent to the disc.
@@
Let $\mathcal{H}_{D}$ denote the horoball in $D(0,1)$ which is the image of the subset in $\mathbf{H}$ consisting of $\tau\in\mathbf{H}$ with $\mathrm{im}(\tau)\geq D$. For $\gamma\in U(1,1)$ such that $\widetilde{\gamma}=\begin{pmatrix}a&b\\ c&d\end{pmatrix}\in\mathrm{PSL}_{2}(\mathbf{R})$, the image $\gamma\mathcal{H}_{D}$ of $\mathcal{H}_{D}$ under $\gamma$ in $D(0,1)$ is a circle tangent to (aic)/(a+ic)(a-ic)/(a+ic) with diameter

### v2

@@ 5.1. Schwarzians and the conformal radius
By Definition 5.1.13, $G_{N}(x)=(F_{N}(x^{1/N}))^{N}$; and by the assumptions $\varphi_{N}(0)=0$ and $\psi_{N}(0)=0$, we obtain the formal identity $\varphi_{N}(z)=\psi_{N}(z^{1/N})^{N}$ (here formally take $z=G_{N}(x)$).
Let $\eta_{1},\eta_{2}$ denote the solutions of the differential equation in Lemma 5.1.8 such that $\eta_{1}(0)=0,\eta^{\prime}_{1}(0)=1,\eta_{2}(0)=1,\eta^{\prime}_{2}(0)=0$; then $\eta_{1},\eta_{2}$ are linearly independent and $\eta_{1}(z)/\eta_{2}(z)=z+O(z^{2})$. Since $\psi_{N}(z)=|F_{N}^{\prime}(0)|^{-1}z+O(z^{2})$, then by Lemma 5.1.8, we have $\psi_{N}=|F_{N}^{\prime}(0)|^{-1}\,\eta_{1}/\eta_{2}$. We deduce
$\varphi_{N}(z)=|F_{N}^{\prime}(0)|^{-N}(\eta_{1}(z^{1/N})/\eta_{2}(z^{1/N}))^{N}.$
Let $\phi_{i}(z)=\eta_{i}(z^{1/N})$. Then
$\phi^{\prime}_{i}(z)=N^{-1}z^{\frac{1}{N}-1}\eta^{\prime}_{i}(z^{1/N})$
and
$\phi^{\prime\prime}_{i}(z)=\frac{1-N}{N^{2}}z^{\frac{1}{N}-2}\eta^{\prime}_{i}+N^{-2}z^{\frac{2}{N}-2}\eta^{\prime\prime}_{i}(z^{1/N}).$
From equation (5.1.9), we have
$4(z-1)^{2}\eta^{\prime\prime}_{i}(z^{1/N})+((N^{2}-1)z^{1-\frac{2}{N}}+z^{2-\frac{2}{N}})\eta_{i}(z^{1/N})=0.$
We rewrite this differential equation in terms of derivatives of $\phi_{i}$ using the above equations and then conclude that $\phi_{1},\phi_{2}$ are solutions to (5.1.15).
@@
The last assertion is just a summary of the above results. ∎
In order to prove Theorem 5.1.4, we need the following formula of the behavior of hypergeometric functions in Lemma 5.1.14 near $z=1$.
Lemma 5.1.18 (See, for instance, [AS92, 15.3.10]).
Given $a\notin\mathbf{Z}_{\leq 0}$, for $|z|<1$, we have
@@
where $\psi$ denotes the logarithmic derivative of the Gamma function. Note that the above hypergeometric function is multivalued around $z=0$, but all different branches are accounted for by the branches of the logarithm. ∎
Proof of Theorem 5.1.4.
Since $F_{N}$ is a covering map of $\mathbf{C}\smallsetminus\mu_{N}$, then local inverse $\psi_{N}$ is naturally defined on $D(0,1)\subset\mathbf{C}\smallsetminus\mu_{N}$. Moreover, since $F_{N}(1)=1$, we have that $\lim_{z\rightarrow 1}\psi_{N}(z)=1$, where $z\in D(0,1)$ approaches $1$. (A priori, we could only conclude that $\lim_{z\rightarrow 1}\psi_{N}(z)$ approaches a cusp of $D(0,1)$ and thus $|\lim_{z\rightarrow 1}\psi_{N}(z)|=1$, which is enough for the rest of the proof; on the other hand, either by the description of the fundamental domain in Lemma 5.2.1 or by the computation of the rest of the proof that $\lim_{z\rightarrow 1}\psi_{N}(z)$ is a positive real number, we indeed have $\lim_{z\rightarrow 1}\psi_{N}(z)=1$ due to our assumption that $\psi_{N}(0)=0$.)
By Lemma 5.1.14, we have $\psi_{N}(z)=|F^{\prime}_{N}(0)|^{-1}s_{N}(z^{N})$. In particular,
$|F^{\prime}_{N}(0)|^{-1}\lim_{z\in D(0,1),z\rightarrow 1}s_{N}(z^{N})=1.$
Thus, by (5.1.17) and (5.1.19), we have
[…]
Basic properties of the Gamma function [AS92, 6.1.18] transform the latter expression into
$\gamma_{N}=2^{4/N}\frac{\displaystyle{\Gamma\left(1+\frac{1}{2N}\right)^{2}\Gamma\left(1-\frac{1}{N}\right)}}{\displaystyle{\Gamma\left(1-\frac{1}{2N}\right)^{2}\Gamma\left(1+\frac{1}{N}\right)}}.$
Then by [AS92, 6.1.33], we also have
$\log\gamma_{N}=\frac{\log 16}{N}+\sum_{k=1}^{\infty}\frac{(2^{2k}-1)}{2^{2k-1}(2k+1)}\cdot\frac{\zeta(2k+1)}{N^{2k+1}}.$
We obtain (5.1.6) by taking the exponential of the above formula. ∎
@@
5.2. Geometry of $\Gamma_{N}$ and a uniform growth estimate of $F_{N}$
Our second aim in the present § 5 is the uniform supremum growth estimate Lemma 5.2.17 of the universal covering map $F_{N}:D(0,1)\to\mathbf{C}\smallsetminus\mu_{N}$ near the boundary, as both the circle $|x|=r$ and the level $N$ vary. This result is subsumed by the more precise bound of Kraus and Roth [KR16, Theorems 1.2 and 1.10]; our treatment is self-contained. In § 6 we will refine this supremum growth bound (uniformly exponential in $\frac{N}{1-r}$) to an integrated growth bound (uniformly linear in $\frac{N}{1-r}$).
The idea of the proof is to use the symmetry of $\mathbf{C}\smallsetminus\mu_{N}$ and the action of the Fuchsian group $\widetilde{\Gamma}_{N}$ to reduce the question to the study of the asymptotic of $\widetilde{F}_{N}$ near the cusp $\tau=i\infty$ and we study this asymptotic using the explicit description of the inverse of $\widetilde{F}_{N}$ given in Lemma 5.1.14.
We begin in this subsection by explicitly describing the Fuchsian group $\widetilde{\Gamma}_{N}$.
Proposition 5.2.1.
The stabilizer of $i\infty$ in $\widetilde{\Gamma}_{N}$ is generated by
$\widetilde{t}_{N}:=\left(\begin{matrix}1&2\cot(\pi/2N)\\ 0&1\end{matrix}\right).$
The group $\widetilde{\Gamma}_{N}$ is the free group on $N$ generators generated by $\widetilde{t}_{N}$ and its conjugates by powers of $\widetilde{r}_{N}$ in (5.1.3). A fundamental domain $\Omega\subset D(0,1)$ for $\Gamma_{N}$ is given by the region with $2N$ cusps given by half-integer powers of $\zeta_{N}=\exp(2\pi i/N)$ and bounded by geodesics connecting adjacent cusps.
Proof.
Since $\widetilde{F}_{N}$ is the universal covering map to $\mathbf{C}\smallsetminus\mu_{N}$, then the Fuchsian group $\widetilde{\Gamma}_{N}$ is generated by the stabilizers of the cusps $c$ with $\widetilde{F}_{N}(c)\in\mu_{N}$. If we denote the generator of the stabilizer of $i\infty$ by
(5.2.2) $\widetilde{t}:=\left(\begin{matrix}1&c_{N}\\ 0&1\end{matrix}\right),$
then the stabilizers of the other cusps associated to $\mu_{N}$ are generated by the conjugates of $\widetilde{t}_{N}$ by $\widetilde{r}_{N}$ since $\zeta_{N}^{k}\widetilde{F}_{N}(\tau)=\widetilde{F}_{N}(\widetilde{r}_{N}^{k}\cdot\tau)$ by Lemma 5.1.2.
Consider the Dirichlet domain $\Omega_{N}\subset D(0,1)$ associated to $\Gamma_{N}$ around $x=0$. More precisely, we can describe $\Omega_{N}$ as the region:
$\{z\in D(0,1)\mid d(gx,0)\geq d(x,0),\quad g,g^{-1}\in\{r_{N}^{k}tr_{N}^{-k}\},k=0,1,\ldots,N-1\}$
The region in $D(0,1)$ such that $d(gx,0)\geq d(x,0)$ and $d(g^{-1}x,0)\geq d(x,0)$ for $g=r_{N}^{k}tr_{N}^{-k}$ is the region bounded by two geodesics starting at $\zeta_{N}^{k}$ going in opposite directions and intersecting the boundary at $\zeta_{N}^{k}e^{\pm i\theta}$ where $c_{N}=2\cot(\theta/2)$. There are exactly $2N$ such arcs corresponding to the $N$ generators and their inverses. In particular, if $\theta<\pi/N$ is too small, the fundamental region will have infinite volume, whereas if $\theta>\pi/N$ is too big, then the Dirichlet domain will only contain at most $N$ cusps. Since $\mathbf{H}/\widetilde{\Gamma}_{N}$ has $N+1$ cusps and $\widetilde{\Gamma}_{N}$ has finite covolume, we must have $c_{N}=2\cot(\pi/2N)$ and these geodesics intersecting at $\zeta_{N}^{k+1/2}$, $k=0,\ldots,N-1$. ∎
Now we consider the group associated to $\widetilde{F}_{N}^{N}$.
Definition 5.2.3.
Let $\widetilde{\Phi}_{N}$ denote the group $\langle\widetilde{\Gamma}_{N},\widetilde{r}_{N}\rangle=\langle\widetilde{r}_{N},\widetilde{t}_{N}\rangle$ and let $\Phi_{N}$ denote the corresponding lattice in $\mathrm{PSU}(1,1)$.
Corollary 5.2.4.
The function $\widetilde{F}^{N}_{N}$ is invariant under $\widetilde{\Phi}_{N}$ and $\widetilde{\Phi}_{N}$ is the largest subgroup of $\mathrm{PSL}_{2}(\mathbf{R})$ with this property. A fundamental domain $\Omega^{\prime}_{N}$ for $\Phi_{N}$ in $D(0,1)$ is given by the hyperbolic quadrilateral with vertices $0,\zeta_{N}^{-1/2},1,\zeta_{N}^{1/2}$. Translated to $\mathbf{H}$ this is bounded by geodesics from $i$ to $\cot(\pi/2N)$ to $i\infty$ to $\cot(\pi/2N)$ and back to $i$.
Proof.
The statements follow directly from Lemma 5.1.2 and Proposition 5.2.1. ∎
@@ 5.2. Geometry of $\Gamma_{N}$ and a uniform growth estimate of $F_{N}$
Let $\widetilde{s}\in\mathrm{PSL}_{2}(\mathbf{R})$ be a rotation of order $2N$ such that $\widetilde{s}^{2}=\widetilde{r}_{N}$. Then
(5.2.6) $1-\widetilde{F}^{N}_{N}(\widetilde{s}\cdot\tau)=\frac{1}{1-\widetilde{F}^{N}_{N}(\tau)}.$
@@
Recall our normalization of $\widetilde{F}_{N}$ that $\widetilde{F}_{N}(i)=0$ and $\widetilde{F}_{N}(i\infty)=1$. By Corollary 5.2.4, $\tau=\pm\cot(\pi/2N)$ (in the same $\widetilde{\Phi}_{N}$-orbit) is the other cusp of $\widetilde{\Phi}_{N}$ and thus $\widetilde{F}_{N}(\pm\cot(\pi/2N))=\infty$. Moreover, both sides of (5.2.6) are uniformizers of $\mathbf{H}/\widetilde{\Phi}_{N}$ which take the value 0 at the cusp $\cot(\pi/2N)$ and $\infty$ at the cusp $i\infty$. This specifies them uniquely up to $z\mapsto\lambda z$ scalings. However, this last ambiguity is removed by noting that both sides are $1$ at $\tau=i$. ∎
Remark 5.2.7.
The group $\widetilde{\Phi}_{N}$ is contained with index two in the larger group $\widetilde{\Psi}_{N}=\langle\widetilde{s},\widetilde{t}_{N}\rangle$. The group $\widetilde{\Psi}_{N}$ has a fundamental domain consisting of the points $0,1,\zeta_{N}^{1/2}$. But this is none other than a hyperbolic triangle with angles $\{\alpha,\beta,\gamma\}=\{\pi/N,0,0\}$, whose conformal mapping from $\mathbf{H}$ is given by Schwarz triangle functions (see for instance [Car54, §404 on p. 185]). This suggests that $\widetilde{F}_{N}$ should directly be related to Schwarz triangle functions, which leads to a direct description of the inverse functions $\psi_{N}$ and $\varphi_{N}$ in terms of hypergeometric functions in Lemma 5.1.14.
In order to study the behavior of $\widetilde{F}_{N}$ near $z=1$, we use the explicit formula for its inverse $\psi_{N}$ given in Lemma 5.1.14. The following lemma gives the asymptotic of the function $s_{N}$ used in formula of $\psi_{N}$.
Lemma 5.2.8.
Fix a real constant $M_{0}>0$. For $M\geq M_{0}$ and $|z|<e^{-MN}$, we have the uniform estimate:
(5.2.9) $\left|\frac{s_{N}(1-z)}{\gamma_{N}}-(1-z)^{1/N}\frac{-\log z+2\gamma-2\psi(1/2+1/2N)}{-\log z+2\gamma-2\psi(1/2-1/2N)}\right|\ll\frac{|z|}{N}$
where the implied estimates depend on $M_{0}$ but not on $N,M$ and $\psi$ is the digamma function as in Lemma 5.1.18.
Proof.
By Lemma 5.1.18, for $a=1/2\pm 1/2N$ and $|z|<1$, we have the following equality:
@@
We first prove that the coefficients in this power series are uniformly bounded. Since $|a|<1$, we have $|(a)_{k}|/k!<1$; moreover by the properties of $\psi$-functions (see for instance [AS92, 6.3.5, 6.3.14]), we have $\psi(z)$ is negative and strictly increasing for $0<z<1$ and $|\psi(k)-\psi(k+a)|>|\psi(k+1)-\psi(k+1+a)|$; thus $|\psi(k)-\psi(k+a)|$ is maximized when $k=1$ and a=1/21/4a=1/2-1/4. This immediately leads to the uniform estimates:
$\left|\sum_{k=1}^{\infty}\frac{(a)_{k}(a)_{k}}{k!^{2}}z^{k}\right|,\quad\left|\sum_{k=1}^{\infty}\frac{(a)_{k}(a)_{k}}{k!^{2}}z^{k}\left(2(\psi(k+1)-\psi(k+a))\right)\right|\ll|z|<e^{-MN}$
For $|z|<e^{-MN}$, we also have $\mathrm{Re}(\log z)\leq-MN$, and so in particular $|\log z|\geq MN$ regardless of the branch of logarithm. Combined with Lemma 5.1.14 and (5.2.10), this leads to the estimate
$\frac{s_{N}(1-z)}{\gamma_{N}(1-z)^{1/N}}=\frac{-\log(z)+2\psi(1)-2\psi(1/2+1/2N)+O(z)}{-\log(z)+2\psi(1)-2\psi(1/2-1/2N)+O(z)}$
@@
Fix a pair of real positive numbers $M_{0}>0$ and $\epsilon>0$. Consider any $M\geq M_{0}$, and let $\widetilde{\Omega}^{\prime}_{N}\subset\mathbf{H}$ denote the fundamental domain for $\widetilde{\Phi}_{N}$ corresponding to $\Omega^{\prime}_{N}$ in Corollary 5.2.4. If $\tau\in\widetilde{\Omega}^{\prime}_{N}$ has
$\|\widetilde{F}_{N}(\tau)^{N}-1\|<e^{-MN},$
then
(5.2.12) $\mathrm{im}(\tau)>\frac{2N^{2}M}{\pi^{2}}(1-\epsilon)$
once $N\gg_{\epsilon,M_{0}}1$, where the implicit coefficient depends only on $M_{0}$ and $\epsilon$.
@@
By Corollary 5.2.4, $\widetilde{\Omega}^{\prime}_{N}$ is a fundamental domain of $\widetilde{F}^{N}_{N}$ and the only cusp where $\widetilde{F}_{N}^{N}=1$ is at $\tau=i\infty$. So it suffices to consider $\widetilde{F}^{N}_{N}$ in a neighbourhood of the cusp $i\infty$.
For $N$ sufficiently large, the inequality $\|\widetilde{F}_{N}(\tau)^{N}-1\|<e^{-MN}$ implies that $|\widetilde{F}_{N}(\tau)-1|<e^{-M(1-\varepsilon_{0})N}$ for some $\varepsilon_{0}$ that tends to zero as $N$ increases. Recall from the proof of Theorem 5.1.4 that $\psi_{N}(1)=1$ and hence $\widetilde{\psi}_{N}(1)=i\infty$; then it suffices to bound the imaginary part of
$\tau=\widetilde{\psi}_{N}(1-z),\qquad\text{for }|z|<e^{-M(1-\varepsilon_{0})N}.$
@@
Write $(1-z)^{N}=1-Z$, then the same estimate as above implies $|Z|<e^{-M(1-2\varepsilon_{0})N}$ for sufficiently large $N$. Thus we reduce the lemma to the estimate of
(5.2.14) $\tau=i\cdot\frac{\gamma_{N}+s_{N}(1-Z)}{\gamma_{N}-s_{N}(1-Z)},\quad|Z|<e^{-M(1-2\varepsilon_{0})N}.$
Now by Lemma 5.2.8 and [AS92, 6.3.7], we have
$\tau=\frac{i\cot(\pi/2N)}{\pi}\left(2\gamma-\log Z-\psi(1/2-1/2N)-\psi(1/2+1/2N)\right)+O(1),$
where the implicit constant only depends on $M_{0}$ and $\varepsilon_{0}$. The imaginary part of this does not depend on the choice of branch of $\log Z$ and indeed only depends on $|Z|$, and we deduce with this approximation that
$\mathrm{im}(\tau)\geq\frac{\cot(\pi/2N)}{\pi}\left(2\gamma+NM(1-2\varepsilon_{0})-\psi(1/2-1/2N)-\psi(1/2+1/2N)\right)+O(1).$
If we choose $\varepsilon_{0}:=\epsilon/2$, then for $N\gg_{\epsilon,M_{0}}1$ this lower bound clearly exceeds the requisite
$\frac{2N^{2}M(1-\epsilon)}{\pi^{2}}.$
The region (5.2.12) is a horoball for the cusp $i\infty$ in the upper half plane model $\mathbf{H}$ of the hyperbolic plane. Recall that in the Poincaré disc model $D(0,1)$, the horoballs are the euclidean discs inside $D(0,1)$ which are tangent to the boundary circle. The following easy lemma describes how these horoballs transform under the hyperbolic isometry group.
@@
Let $\mathcal{H}_{D}$ denote the image in the Poincaré disc model $D(0,1)$ of the horoball
$\{\tau\in\mathbf{H}\,:\,\mathrm{im}(\tau)\geq D\}$
in the upper half plane model $\mathbf{H}$. For $\gamma\in\mathrm{PSU}(1,1)$ with image $\widetilde{\gamma}=\begin{pmatrix}a&b\\ c&d\end{pmatrix}\in\mathrm{PSL}_{2}(\mathbf{R})$, the image $\gamma\mathcal{H}_{D}$ of $\mathcal{H}_{D}$ under $\gamma$ in $D(0,1)$ is the disc with diameter
@@
tangent to the boundary circle $\mathbf{T}$ at the point (aic)/(a+ic)(a-ic)/(a+ic). ∎
We close this section by using Lemma 5.2.11 to derive a coarse yet fairly uniform upper bound on $\sup_{|x|=r}\log|F_{N}(x)|$. Although not best-possible, it is enough as an input for the logarithmic error term in the Nevanlinna theory estimate in §6.